OpenStudy (anonymous):

Question: The Difference of one number and four times a second number is 32. Find the minimum Product. How far I got: P = x * Y // x-4y=32 solve for one variable in (x-4y=32), then input the found variable into p = x * y. 32+4y = x => P = (32+4y)y = > P= 32y+4y^2 then get the derivative of P, dp/dx= 32+8y ? I'm not sure if I'm doing it right or not, can i get some guidance! Thanks

OpenStudy (anonymous):

\[x-4y=32\] \[P=xy\] \[x=32+4y\] \[P(y)=(32+4y)y=32y+4y^2\]

OpenStudy (anonymous):

looks good

OpenStudy (anonymous):

\[P'(y)=32+8y\] set equal to zero and get \(y=-4\)

OpenStudy (anonymous):

Thank you man!